OpenStudy (anonymous):

What is the slope-intercept form of the equation of a line that passes through (5, -4) and has a slope of 3/4?

OpenStudy (anonymous):

So are you looking for the y=mx + b form?

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\] replace \(x_1\) by 5, \(y_1\) by \(-4\) and \(m\) by \(\frac{3}{4}\)

OpenStudy (anonymous):

ok